Im not sure how to type in this problem I need help so bare with me 3^the sqrt of x^2=4 I need help in how to solve this problem
Answer by stanbon(75887)

3sqrt(x^2) = 4
Divide both sides by 3 to get:
sqrt(x^2)= (4/3)
Square both sides to get:
x^2 = (4/3)^2
x=(4/3) or x=(-4/3)
==============
Cheers,
Stan H.